4. Given the consumption C = 20 + 0.4Yd, and


T = 600 I = 500, G = 600, then;


i. Write the equation of aggregate income and simplify.


ii. Calculate marginal propensity to save and interpret the result.


iii. Find autonomous tax multiplier.


iv. Calculate the change of income when G decrease by 100.


v. Calculate the new equilibrium income when tax equation is given as T = 0.25Y


vi. Calculate the new equilibrium income when I increase to 600.



1
Expert's answer
2022-03-07T10:43:48-0500

1) Y=C+I+G

Y=20+0.4*Y+500+600

Y-0.4Y=1120

0.6Y=1120

Y=1120/0.6=1867.67

2) MPS=1-MPC=1-0.4=0.6

3)mt=MPC/MPS=0.4/0.6=0.67

4)Change of Y=Change of G*mg

mg=1/MPS=1/0.6=1.67

Change of Y=-100*1.67=-167

5) C=C0+MPC*(Y-T)

CO+0.4(Y-600)=20+0.4Y

CO+0.4Y-240=20+0.4Y

C0-240=20

C0=260

New function of consumption: C=260+0.4(Y-0.25Y)

C=260+0.3Y

Y=260+0.3Y+500+600

0.7Y=1360

Y=1360/0.7=1942.86

6) Y=20+0.4Y+600+600

0.6Y=1220

Y=1220/0.6=2033.33